# HoPE: A Novel Positional Encoding Without Long-Term Decay for Enhanced Context Awareness and Extrapolation

### v1

@@ 5.2.1 Evaluation
Perplexity (PPL) is the most commonly used metric for evaluating a model’s extrapolation capability. We conduct our evaluation on a subset of the C4 dataset Raffel et al. (2019a) with 1,000 samples by comparing the zero-shot perplexity of the last 256 tokens across different input lengths.
Copy Task
@@
The copy capability is one of the most fundamental abilities of language models and is closely related to token order. Many previous works Liu et al. (2023); Golovneva et al. (2024); Lv et al. (2024) on model structure optimization have designed similar tasks to evaluate the models’ effectiveness. Based on these studies, We designed our copy task. Specifically, we constructed a test set containing 500 samples, with each sample consisting of multiple sequences. Each sequence has an average length of 12 tokens, with a unique 8-gram prefix and 4-gram suffix. During testing, we concatenate a specific number of sequences with the prefix of a certain $i$-th sequence (queried sequence) to serve as the model’s input. The model’s objective is to output the suffix of the queried prefix, with the middle sequence selected as the queried one. Figure 3 illustrates a specific example input case.
Figure 3: Specific input example of copy task.
Few-shot Task
@@
Few-shot learning is another core ability of the model and serves as the foundation for instruction following. We created a test set with 600 samples selected from three tasks (SST-2, QNLI and RTE) in the GLUE Wang et al. (2018) benchmark. For each input, we concatenate few-shot examples, a set of meaningless sentences, and the queried input. Specific examples can be found in Figure 8. As for the evaluation metric, instead of focusing on actual accuracy, we emphasize whether the model’s output falls into the label sets from the few-shot examples. For instance, if the label set in the contextual examples is ${0,1}$, the model’s output, whether 0 or 1, will be counted. And we define this as a measure of follow ability (FA). We present the average performance across the three tasks in the main text, and more detailed results for each task please refer to the Appendix C.
@@
We set the rotary base $b=10,000$ in HoPE. Other settings are consistent with those in Section 3. We evaluate the proposed HoPE against a range of established baselines, including RoPE Su et al. (2021), ALiBi Press et al. (2021), KERPLE Chi et al. (2022a), and FIRE Li et al. (2024), as well as two typical RoPE-based extrapolation methods: PI Chen et al. (2023) and YaRN Peng et al. (2023).
5.2.2 Result
@@
The results for perplexity (PPL), copy task, and few-shot task are presented in Figure 4, Table 1 and Table 2, respectively. From these results, we can draw the following conclusions:
@@
(1) From all perspectives in the figure and tables above, it can be confirmed that our approach significantly enhances the context awareness and extrapolation of the original RoPE. As shown in Figure 4, our method noticeably smooths the increase in PPL observed in RoPE, achieving low PPL even with training lengths 4 times longer or more. It records a PPL of 8.5241 at 512, and 13.0257 at 4,096. Table 1 and 2 further demonstrate that that our approach not only improves extrapolation but also enhances context awareness within the training length. Specifically, compared to RoPE, the model’s copy ability increased from an average of 23.80 to 60.23, while its few-shot capability improved from an average of 54.10 to 79.20.
@@
(2) When combined with extrapolation methods like PI and YaRN, our method achieves even better extrapolation results. As shown in Table 1 and 2, our method, when integrated with YaRN, achieves the best overall performance across all PEs, with an average score of 68.63 in the copy task and 93.77 in the few-shot task. However, as also noted in tables, while both PI and YaRN enhance extrapolation, they appear to negatively impact the model’s context awareness within the training length.
@@
(3) Relying solely on perplexity (PPL) to measure extrapolation is not reliable. In some cases, the PPL measurements (shown in Figure 4) contradict the performance in other tasks (shown in Table 1 and 2), indicating that PPL may not prove a method’s ability to effectively utilize global information. It might reflect “pseudo” extrapolation, as seen in the result of ALiBi and KERPLE. From Table 1 and 2, it is evident that ALiBi’s actual extrapolation is poor, and while KERPLE shows some extrapolation, it is not as strong as the PPL suggests and performs slightly worse in few-shot performance.
@@
Table 2: The results of the few-shot experiment. We measure the model’s following ability (FA), which counts the instances when the output includes one of the label sets from the examples. The leading results are highlighted with bold fonts.
@@ 6 Ablation Study of HoPE
Result
As shown in Table 3, removing the “activated” components (AB1) results in a significant improvement in both context awareness and extrapolation, with an average increase of 14.5 points compared to RoPE. This outcome suggests that these “activated” components indeed contribute to the model’s short-cut learning, hindering optimal learning.
@@
7 Effect Scaling on Model Size and Training Length of HoPE
In this section, we explore the scaling effects of our method with respect to model size and training length.
7.1 Scaling on Model Size
Set up
We increased the scale of model from 125M to 300M and compared the performance of all position encodings in terms of perplexity (PPL) and the copy task. The details of model configurations can be found in Appendix A.
Figure 5: Perplexity Comparison with the model size 300M.
Method Sequence Number
30 40 50 60 70 80 Avg.
ALiBi 81.30 68.80 58.60 33.20 17.80 12.50 45.37
KERPLE 84.80 78.20 73.20 68.80 47.40 36.90 64.88
FIRE 81.80 77.20 58.40 25.00 3.40 0.60 41.07
RoPE 87.80 85.40 55.80 0.40 0.00 0.00 38.23
+PI 79.20 67.80 65.40 22.40 0.60 0.00 39.23
+YaRN 75.80 67.00 65.10 52.80 48.70 49.30 59.78
Our HoPE 89.20 87.40 80.00 57.10 44.70 37.60 66.00
+PI 87.20 84.60 77.20 56.90 48.30 43.40 66.27
+YaRN 88.00 84.60 84.10 61.80 52.00 42.70 68.87
Table 4: Results of the copy task with the model size 300M.
Result
Results are shown in Figure 5 and Table 4. As the model size grows, all methods show improvements in their metrics. Our HoPE’s PPL at a length of 512 decreases from 8.5241 to 7.4629, and at 4192, it drops from 13.0257 to 10.6933. In terms of copy ability, HoPE’s average score increases from 60.23 to 66.00. Additionally, our HoPE consistently demonstrates superior performance, highlighting its robust improvements. Our HoPE yields an average increase of 27.77 points compared to RoPE in the copy task. As for PI and YaRN, although they still show great effectiveness in extrapolation, their performance within the training length declines more noticeably, particularly at sequence number 40, where they show drops of 17.6 and 18.4 points, respectively, compared to RoPE.
7.2 Scaling on Training Length
Set up
We varied the training lengths from 512 to 2048 in a doubling increase. We use PPL and few-shot tasks for evaluation.
Result
The results in Table 5 and Table 6 demonstrate that our approach consistently shows strong performance as the training length increases. The model maintains low PPL even beyond four times the training length, indicating robust extrapolation. Moreover, as the training length extends, the model demonstrates notable improvements in context awareness. As shown in the Table 5 and Table 6, when the training length increases from 512 to 2048, PPL for the test length of 512 decreases from 8.5241 to 7.8704, while performance in the few-shot task improves from 96 to 100.
Train/Test 512 1024 2048 4096 8192 16384
512 8.5241 9.1419 10.1830 13.0257 19.6137 34.8469
1024 8.1137 7.8897 8.1023 8.8615 11.3563 19.0351
2048 7.8704 7.5870 7.3026 7.8704 9.3932 10.1004
Table 5: Perplexity across different training and test lengths. We use underline to highlight the extrapolation results.
Train/Test 512 1024 1536 2048 3072 4096
512 96.00 90.00 88.67 83.00 70.83 67.83
1024 100.00 95.00 96.17 87.67 83.83 81.67
2048 100.00 100.00 99.00 96.00 91.17 86.33
Table 6: Results of Few-Shot Tasks Across Varying Training and Test Lengths.
@@ 8 Conclusion
In this paper, we explore the empirical attention patterns of various positional encodings and observe that position-related attention tend to form a U-shape pattern, benefiting more from local decay rather than global. Our further analysis of RoPE reveals a strong correlation between the U-shape pattern and its learned components. We identify that certain "activated" components and top low-frequency components in RoPE hinder the model’s optimal learning process, limiting its context awareness and extrapolation. Consequently, we propose our method, HoPE, which breaks the principle of long-term decay in theory, allowing for optimal utilization of components for positional encoding. Extensive experiments demonstrate its effectiveness in enhancing both context awareness and extrapolation. We hope our method will inspire the community to examine many of the “taken for granted” design choices in transformers, with the aim of advancing the development of more powerful large language models (LLMs).
@@ Appendix A Model Configuration
Detail settings across model sizes are depicted in Table 7. Other hyperparameters are as follows: the AdamW optimizer is used with $(\beta_{1},\beta_{2})=(0.9,0.999)$, a learning rate of $3e^{-4}$, $2,000$ warm-up steps, and a gradient clipping value of 1. All experiments are conducted on 8 A100 GPUs with a batch size of 64 per GPU.
125M 300M
Training sequence length 512 512
Batch size $32\times 8$ $32\times 8$
Number of iterations 50k 50k
Dropout prob. 0.0 0.0
Attention dropout prob. 0.0 0.0
@@
Feature dimension 768 1024
Intermediate dimension 2688 3584
@@ Appendix B Supplementary Results on the Exploration of Component Effects in RoPE Attention Patterns
We present supplementary experiments with training lengths of 1024, depicted in Figure 6 and  7. We reached the same conclusion as in the main text. These “activated” components still exhibit a U-shaped curve, similar to the final pattern and lower frequency components continue to learn a constant pattern. It can be seen that “activated” and lower frequency components shift further back as the training length increases.
@@
Appendix C Input Examples of Few-shot Tasks
Specific example of few-shot tasks can be found in Figure 8. In practice, We provide 5-shot as training example and dynamically pad with a different number of meaningless sentences to ensure varying output lengths.
Figure 8: Input examples of few-shot tasks. We take SST-2 as an instance.

### v2

@@ 4.2.1 Evaluation
Perplexity (PPL) is a commonly used metric for evaluating a model’s extrapolation capability. We conduct our evaluation on a subset of the C4 dataset Raffel et al. (2019a) with 1,000 samples by comparing the zero-shot perplexity of the last 256 tokens across different input lengths.
In-Context Copying
@@
The copying capability is one of the most fundamental abilities of language models and is closely related to token order. Many previous works Liu et al. (2023); Golovneva et al. (2024); Lv et al. (2024) on model structure optimization have designed similar tasks to evaluate the models’ effectiveness. Based on these studies, We designed our copying task. Specifically, we constructed a test set containing 500 samples, with each sample consisting of multiple sequences. Each sequence has an average length of 12 tokens, with a unique 8-gram prefix and 4-gram suffix. During testing, we concatenate a specific number of sequences with the prefix of a certain $i$-th sequence (queried sequence) to serve as the model’s input. The model’s objective is to output the suffix of the queried prefix, with the middle sequence selected as the queried one. Figure 3 illustrates a specific example input case.
Figure 3: Specific input example of copying task.
Few-shot Following
@@
Few-shot learning is another core ability of the model and serves as the foundation for instruction following. We created a test set with 600 samples selected from three tasks (SST-2, QNLI and RTE) in the GLUE Wang et al. (2018) benchmark. For each input, we concatenate few-shot examples, a set of meaningless sentences, and the queried input. Specific examples can be found in Figure 8. As for the evaluation metric, instead of focusing on actual accuracy, we emphasize whether the model’s output falls into the label sets from the few-shot examples. For instance, if the label set in the contextual examples is ${0,1}$, the model’s output, whether 0 or 1, will be counted. And we define this as a measure of follow ability (FA). We present the average performance across the three tasks in the main text, and detailed results for each task please refer to Appendix C.2.
@@
We set the rotary base $b=10,000$ in HoPE. Other settings are consistent with those in Section 3.1. We evaluate the proposed HoPE against a range of established baselines, including RoPE Su et al. (2021), ALiBi Press et al. (2021), KERPLE Chi et al. (2022a), and FIRE Li et al. (2024), as well as two typical RoPE-based extrapolation methods: PI Chen et al. (2023) and YaRN Peng et al. (2023).
4.2.2 Results
@@
The results for perplexity (PPL), copying task, and few-shot following task are presented in Figure 4, Table 2 and Table 3, respectively. From these results, we can draw the following conclusions:
@@
(1) From all perspectives in the figure and tables above, it can be confirmed that our approach significantly enhances the context awareness and extrapolation of the original RoPE. As shown in Figure 4, our method noticeably smooths the increase in PPL observed in RoPE, achieving low PPL even with training lengths 4 times longer or more. It records a PPL of 8.5241 at 512, and 13.0257 at 4,096. Table 2 and 3 further demonstrate that our approach not only improves extrapolation but also enhances context awareness within the training length. Specifically, compared to RoPE, the model’s copy ability increased from an average of 23.80 to 60.23, while its few-shot following capability improved from an average of 54.10 to 79.20.
@@
(2) When combined with extrapolation methods like PI and YaRN, our method achieves even better extrapolation results. As shown in Table 2 and 3, our method, when integrated with YaRN, achieves the best overall performance across all PEs, with an average score of 68.63 in the copy task and 93.77 in the few-shot following task. However, as also noted in the tables, while both PI and YaRN enhance extrapolation, they appear to negatively impact the model’s context awareness within the training length.
@@
(3) Relying solely on perplexity (PPL) to measure extrapolation is not reliable. In some cases, the PPL measurements (shown in Figure 4) contradict the performance in other tasks (shown in Table 2 and 3), indicating that PPL may not prove a method’s ability to effectively utilize global information. It might reflect “pseudo” extrapolation, as seen in the result of ALiBi and KERPLE. From Table 2 and 3, it is evident that ALiBi’s actual extrapolation is poor, and while KERPLE shows some extrapolation, it is not as strong as the PPL suggests and performs slightly worse in few-shot following performance.
@@
Table 3: The results of the few-shot following experiment. We measure the model’s following ability (FA), which counts the instances when the output includes one of the label sets from the examples. The leading results are highlighted with bold fonts.
4.3 Attention Patterns in HoPE
Figure 5: Attention patterns in HoPE and RoPE, both within (top) and beyond (bottom) training length.
We provide a observation about the position-related attention patterns of HoPE in Figure 5. As shown in the upper subplot, HoPE demonstrates a U-shaped fluctuation similar to RoPE within the training length. However, the positional fluctuation in HoPE is somewhat milder, indicating that HoPE may be better adapted to long-context tasks which requires greater focus on semantic information. Regarding the extrapolation patterns (shown in the lower subplot of Figure 5), HoPE does not exhibit the out-of-distribution (OOD) behavior seen in RoPE, which could account for its superior extrapolation capabilities.
@@ 5 Ablation Study of HoPE
Results
As shown in Table 4, removing the “activated” components (AB1) results in a significant improvement in both context awareness and extrapolation, with an average increase of 14.5 points over RoPE. This outcome suggests that these components indeed contribute to the model’s short-cut learning, hindering optimal learning.
@@ Results
6 Scalability of HoPE
Based on our understanding of HoPE’s advantages, derived from a series of empirical experiments with small models and toy tasks, we conducted further comparative experiments on real-world tasks. These experiments involved training large language models from scratch, comparing RoPE with HoPE.
Settings
We trained a 3B Llama-based model with a training length of 8192 over 120k steps, processing approximately 500B tokens in total. The model configurations and training details can be found in Appendix A. For evaluation, we selected 8 general-purpose benchmarks, including MMLU(5-shot) Hendrycks et al. (2020), MMLU-PRO(5-shot) Wang et al. (2024), GPQA(0-shot) Rein et al. (2023), BBH(3-shot) Srivastava et al. (2022), WinoGrande(5-shot) Sakaguchi et al. (2019), GSM8k(8-shot) Cobbe et al. (2021), MATH(4-shot) Lightman et al. (2023) and DROP(3-shot) Dua et al. (2019). These datasets allow for a comprehensive evaluation of the model’s performance across various tasks such as multiple-choice questions, reasoning, and numerical computations. The shot count settings follow several works Dubey et al. (2024); Bai et al. (2023) and we use OpenCompass Fu et al. (2024) to compute the results.
Results
As shown in Table 1, our HoPE achieves better overall performance compared to RoPE, with a higher average score (26.42) and notable improvements on tasks like MMLU, GPQA, and DROP. We also provide the results of each dataset for every 100B token updates in Appendix D, which further illustrate HoPE’s consistent superiority across most datasets and highlight its robust advantages. These findings highlight the potential of HoPE to drive the next generation of state-of-the-art LLMs by replacing the widely used RoPE.
@@ 7 Conclusion
In this paper, we explore the empirical attention patterns of various positional encodings and observe that position-related attention tend to form a U-shape pattern, benefiting more from local decay rather than global. Our further analysis of RoPE reveals a strong correlation between the U-shape pattern and its learned components. We identify that certain "activated" components and top low-frequency components in RoPE hinder the model’s optimal learning process, limiting its context awareness and extrapolation. Consequently, we propose our method, HoPE, which breaks the principle of long-term decay in theory, allowing for optimal utilization of components for positional encoding. Extensive experiments demonstrate its effectiveness in enhancing both context awareness and extrapolation.
@@ Appendix A Model Configuration
Detail settings across model sizes are depicted in Table 5. Other hyperparameters are as follows: the AdamW optimizer is used with $(\beta_{1},\beta_{2})=(0.9,0.999)$, a learning rate of $3e^{-4}$, $2,000$ warm-up steps, and a gradient clipping value of 1. Experiments for the 125M models are conducted on 8 A100 GPUs, while those for the 3B models use 256 A100 GPUs.
Hyperparameters 125M 3B
Training sequence length 512 8192
Batch size $64\times 8$ $2\times 256$
Number of Iterations 50k 120K
Number of Layers 12 34
@@
Feature Dimension 768 2048
Intermediate Dimension 2688 8704
@@ Appendix B Supplementary Results on the Exploration of Component Effects in RoPE Attention Patterns
We present supplementary experiments with training lengths of 1024, depicted in Figure 6 and  7. We reached the same conclusion as in the main text. It can be seen that “activated” and lower frequency components shift further back as the training length increases. These “activated” components still exhibit a U-shaped curve, similar to the final pattern. The lower frequency components continue to learn a constant pattern.
@@
Appendix C Supplementary for Few-shot Following Tasks
C.1 Input Examples of Few-shot Following Tasks
Specific example of few-shot following tasks can be found in Figure 8. In practice, We provide 5-shot training examples as context. For each instance, we dynamically pad it with a different number of meaningless sentences to ensure the various input lengths.
Figure 8: Input examples of few-shot following tasks. We take SST-2 as an instance.
C.2 Detail Results on Few-shot Following Task
We present the average results in Table 3 of main text. Detail results in three tasks are depicted in Table 6.
Method SST-2 QNLI RTE
256 512 768 1024 1280 Avg. 256 512 768 1024 1280 Avg. 256 512 768 1024 1280 Avg.
ALiBi 99.00 58.00 29.00 4.00 0.00 38.00 100.00 99.00 79.00 0.00 0.00 55.60 100.00 100.00 98.00 15.00 3.00 63.20
KERPLE 50.50 37.50 18.00 9.00 6.00 24.20 100.00 95.00 32.00 7.00 4.00 47.60 81.00 78.00 17.00 28.00 39.00 48.60
FIRE 99.50 99.00 76.50 31.50 3.00 61.90 99.00 99.00 18.00 19.00 12.00 49.40 63.00 61.00 2.00 8.00 43.00 35.40
RoPE 99.50 96.00 20.00 0.00 0.00 43.10 95.00 95.00 85.00 22.00 9.00 61.20 100.00 100.00 49.00 29.00 12.00 58.00
+ PI 99.50 96.50 33.00 4.50 0.00 46.70 100.00 100.00 94.00 75.00 11.00 76.00 95.00 100.00 73.00 36.00 13.00 63.40
+ YaRN 99.50 57.00 73.00 73.00 0.00 60.50 100.00 100.00 100.00 95.00 9.00 80.80 100.00 100.00 100.00 77.00 52.00 85.80
Our HoPE 99.00 95.00 45.50 52.50 32.00 64.80 100.00 100.00 100.00 88.00 37.00 85.00 100.00 100.00 78.00 63.00 98.00 87.80
+ PI 97.50 97.50 97.00 37.00 89.00 83.60 100.00 100.00 100.00 100.00 55.00 91.00 100.00 98.00 54.00 99.00 11.00 72.40
+ YaRN 100.00 97.50 76.00 69.00 98.00 88.10 92.00 80.00 99.00 99.00 100.00 94.00 100.00 99.00 100.00 98.00 99.00 99.20
Table 6: Detail results on three few-shot tasks.
Figure 9: Performance comparison across 8 benchmarks using a 3B Llama-based model with different number of tokens trained.
Appendix D Supplementary Results of the 3B Llama-based Model
In this section, we provide the supplementary results of the 3B Llama-based Model with a training length 8192. As depicted in Figure 9, we evaluate the results on each dataset every 100B token updates.